# Data-Driven Logistic Regression Ensembles With Applications in Genomics

### v5

@@ 8.11 Results
Tables 38-51 contain the full results for the gene expression data benchmark study in Section 7 of the article. For each data set we report the performance of the eleven methods listed in Section 2 of the article for the considered training set proportions and the different numbers of genes $p$ preserved after the pre-processing step, where we considered $p=$ 100, 250, 500 and 1,000. To simplify the comparison for each setting, we report relative performance of the methods in terms of MR and TL. That is, the performance reported for each method is relative to the best performing method, i.e. the ratio of its performance to the best performing method. Thus, the best performing method in each column of the tables corresponds the value 1.00.
Table 9: Mean misclassification rate, sensitivity and specificity for Scenario 1, $\rho=$ 0.2, $n=$ 50, $p=$ 1,000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$ $\mathbf{\pi_{1}}$ Method MR SE SP MR SE SP MR SE SP Lasso 0.29 0.53 0.84 0.20 0.68 0.87 0.17 0.74 0.89 Elastic Net 0.27 0.56 0.85 0.18 0.72 0.89 0.14 0.78 0.91 Adaptive 0.35 0.33 0.88 0.26 0.49 0.91 0.22 0.59 0.90 Relaxed 0.29 0.52 0.83 0.21 0.68 0.86 0.18 0.74 0.88 MCP 0.33 0.46 0.82 0.27 0.53 0.85 0.25 0.60 0.85 0.4 SIS-SCAD 0.33 0.50 0.78 0.30 0.56 0.80 0.28 0.61 0.79 Split-Lasso-10 0.24 0.61 0.86 0.15 0.77 0.91 0.10 0.84 0.94 Split-EN-10 0.23 0.62 0.87 0.15 0.77 0.91 0.10 0.85 0.94 RGLM-100 0.23 0.63 0.86 0.16 0.70 0.92 0.13 0.76 0.94 RF-500 0.24 0.56 0.89 0.17 0.66 0.95 0.13 0.73 0.96 XGB 0.34 0.52 0.76 0.31 0.56 0.77 0.28 0.59 0.79 $\mathbf{\pi_{1}}$ Method MR SE SP MR SE SP MR SE SP Lasso 0.25 0.42 0.91 0.19 0.56 0.92 0.17 0.55 0.95 Elastic Net 0.24 0.42 0.92 0.17 0.60 0.93 0.14 0.62 0.96 Adaptive 0.29 0.13 0.97 0.24 0.29 0.96 0.22 0.29 0.98 Relaxed 0.26 0.41 0.89 0.19 0.57 0.91 0.17 0.58 0.93 MCP 0.28 0.31 0.90 0.25 0.40 0.90 0.24 0.34 0.94 0.3 SIS-SCAD 0.29 0.36 0.87 0.26 0.42 0.88 0.25 0.38 0.90 Split-Lasso-10 0.22 0.48 0.92 0.14 0.67 0.95 0.10 0.71 0.97 Split-EN-10 0.21 0.50 0.92 0.13 0.67 0.95 0.10 0.73 0.97 RGLM-100 0.21 0.49 0.93 0.16 0.55 0.97 0.15 0.52 0.99 RF-500 0.23 0.37 0.96 0.17 0.46 0.98 0.17 0.43 0.99 XGB 0.30 0.41 0.83 0.27 0.45 0.85 0.26 0.44 0.86 $\mathbf{\pi_{1}}$ Method MR SE SP MR SE SP MR SE SP Lasso 0.19 0.25 0.96 0.15 0.39 0.96 0.14 0.39 0.97 Elastic Net 0.19 0.25 0.97 0.14 0.44 0.97 0.12 0.48 0.98 Adaptive 0.21 0.05 0.99 0.19 0.11 0.99 0.18 0.14 0.99 Relaxed 0.19 0.25 0.96 0.16 0.41 0.95 0.14 0.46 0.95 MCP 0.21 0.14 0.96 0.19 0.21 0.96 0.19 0.20 0.96 0.2 SIS-SCAD 0.22 0.14 0.96 0.20 0.20 0.95 0.19 0.20 0.96 Split-Lasso-10 0.17 0.30 0.97 0.12 0.53 0.97 0.09 0.58 0.99 Split-EN-10 0.17 0.31 0.97 0.11 0.55 0.97 0.09 0.61 0.99 RGLM-100 0.17 0.27 0.98 0.14 0.34 0.99 0.14 0.31 1.00 RF-500 0.18 0.18 0.99 0.16 0.25 1.00 0.15 0.21 1.00 XGB 0.23 0.27 0.91 0.20 0.32 0.92 0.20 0.33 0.91
Table 10: Mean misclassification rate, sensitivity and specificity for Scenario 1, $\rho=$ 0.5, $n=$ 50, $p=$ 1,000.
$\mathbf{\boldsymbol{\zeta}=0.1}$ $\mathbf{\boldsymbol{\zeta}=0.2}$ $\mathbf{\boldsymbol{\zeta}=0.4}$ $\mathbf{\pi_{1}}$ Method MR SE SP MR SE SP MR SE SP Lasso 0.16 0.76 0.89 0.12 0.84 0.91 0.10 0.86 0.93 Elastic Net 0.15 0.78 0.89 0.10 0.86 0.92 0.08 0.89 0.94 Adaptive 0.20 0.65 0.91 0.14 0.78 0.92 0.11 0.83 0.93 Relaxed 0.17 0.77 0.87 0.13 0.84 0.90 0.11 0.87 0.91 MCP 0.21 0.69 0.86 0.17 0.76 0.87 0.16 0.76 0.89 0.4 SIS-SCAD 0.22 0.66 0.85 0.20 0.72 0.85 0.19 0.72 0.87 Split-Lasso-10 0.14 0.80 0.89 0.08 0.88 0.93 0.06 0.92 0.96 Split-EN-10 0.14 0.81 0.89 0.08 0.89 0.94 0.05 0.92 0.96 RGLM-100 0.14 0.79 0.90 0.09 0.86 0.94 0.07 0.88 0.96 RF-500 0.14 0.78 0.91 0.09 0.85 0.95 0.07 0.87 0.98 XGB 0.23 0.66 0.84 0.20 0.72 0.85 0.19 0.71 0.87 $\mathbf{\pi_{1}}$ Method MR SE SP MR SE SP MR SE SP Lasso 0.15 0.64 0.94 0.11 0.75 0.95 0.09 0.80 0.95 Elastic Net 0.14 0.67 0.94 0.10 0.78 0.96 0.08 0.84 0.96 Adaptive 0.20 0.40 0.97 0.15 0.58 0.97 0.11 0.71 0.96 Relaxed 0.16 0.64 0.92 0.12 0.75 0.93 0.10 0.82 0.93 MCP 0.21 0.50 0.92 0.17 0.58 0.93 0.16 0.66 0.91 0.3 SIS-SCAD 0.22 0.44 0.93 0.19 0.51 0.93 0.18 0.57 0.92 Split-Lasso-10 0.13 0.70 0.94 0.08 0.82 0.96 0.06 0.88 0.97 Split-EN-10 0.13 0.71 0.94 0.08 0.83 0.96 0.05 0.89 0.97 RGLM-100 0.13 0.68 0.95 0.09 0.75 0.97 0.07 0.80 0.98 RF-500 0.13 0.65 0.96 0.10 0.71 0.98 0.08 0.77 0.99 XGB 0.21 0.56 0.89 0.18 0.59 0.91 0.19 0.61 0.90 $\mathbf{\pi_{1}}$ Method MR SE SP MR SE SP MR SE SP Lasso 0.13 0.55 0.95 0.09 0.70 0.96 0.08 0.70 0.97 Elastic Net 0.12 0.59 0.96 0.08 0.74 0.97 0.07 0.75 0.98 Adaptive 0.16 0.30 0.98 0.12 0.48 0.98 0.11 0.51 0.98 Relaxed 0.13 0.57 0.95 0.10 0.71 0.95 0.09 0.72 0.96 MCP 0.17 0.35 0.95 0.16 0.44 0.95 0.16 0.38 0.96 0.2 SIS-SCAD 0.18 0.27 0.97 0.16 0.36 0.96 0.16 0.32 0.97 Split-Lasso-10 0.11 0.63 0.96 0.07 0.80 0.97 0.05 0.81 0.98 Split-EN-10 0.11 0.63 0.96 0.06 0.80 0.97 0.05 0.82 0.98 RGLM-100 0.11 0.57 0.97 0.08 0.69 0.98 0.07 0.65 0.99 RF-500 0.12 0.52 0.98 0.08 0.64 0.99 0.08 0.58 1.00 XGB 0.18 0.48 0.90 0.16 0.54 0.91 0.15 0.51 0.93
Table 11: Mean misclassification rate, sensitivity and specificity for Scenario 1, $\rho=$ 0.8, $n=$ 50, $p=$ 1,000.
[…]

### v6

@@ 8.11 Results
Tables 62-75 contain the full results for the gene expression data benchmark study in Section 6 of the article. For each data set we report the performance of the fourteen methods listed in Section 3 of the article for the considered training set proportions and the different numbers of genes $p$ preserved after the pre-processing step, where we considered $p=$ 100, 250, 500 and 1,000. To simplify the comparison for each setting, we report relative performance of the methods in terms of ACC and TSL. That is, the performance reported for each method is relative to the best performing method, i.e. the ratio of its performance to the best performing method. Thus, the best performing method in each column of the tables corresponds the value 1.00.
Table 13: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.2, $n=$ 50, $p=$ 1000.
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.73 0.58 0.84 0.81 1.11 0.87 0.81 0.91 0.96 0.57 0.90 0.83 0.94 0.97 0.47
Split-EN-10 0.73 0.58 0.84 0.82 1.09 0.88 0.82 0.91 0.96 0.56 0.90 0.84 0.94 0.97 0.46
Lasso 0.69 0.54 0.81 0.76 1.23 0.81 0.72 0.87 0.90 0.85 0.82 0.72 0.89 0.91 0.77
Elastic Net 0.71 0.55 0.82 0.78 1.17 0.83 0.76 0.89 0.92 0.74 0.86 0.77 0.91 0.94 0.64
Adaptive 0.60 0.20 0.92 0.65 1.32 0.76 0.55 0.89 0.86 1.00 0.78 0.58 0.91 0.88 0.95
Relaxed 0.69 0.54 0.81 0.76 1.28 0.80 0.73 0.85 0.89 1.11 0.82 0.72 0.88 0.90 0.96
0.4 MCP 0.66 0.46 0.81 0.71 1.32 0.74 0.59 0.83 0.82 1.10 0.75 0.58 0.85 0.82 1.06
SIS-SCAD 0.65 0.51 0.75 0.70 1.56 0.71 0.58 0.79 0.78 1.27 0.72 0.59 0.80 0.78 1.31
RuleFit 0.67 0.55 0.76 0.72 1.84 0.73 0.61 0.82 0.80 1.40 0.75 0.60 0.85 0.83 1.20
RE-Lasso-100 0.73 0.60 0.84 0.82 1.10 0.86 0.78 0.92 0.95 0.61 0.88 0.79 0.95 0.96 0.54
RE-EN-100 0.74 0.60 0.84 0.82 1.11 0.87 0.80 0.92 0.96 0.58 0.89 0.81 0.95 0.97 0.50
RGLM-100 0.71 0.46 0.90 0.82 1.15 0.83 0.66 0.95 0.96 0.94 0.85 0.66 0.97 0.97 0.91
RF-500 0.73 0.61 0.83 0.82 1.06 0.85 0.75 0.92 0.95 0.79 0.87 0.75 0.95 0.96 0.76
XGB 0.65 0.53 0.74 0.69 1.29 0.73 0.61 0.81 0.81 1.07 0.74 0.60 0.83 0.81 1.04
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.76 0.44 0.92 0.82 1.01 0.88 0.72 0.95 0.96 0.53 0.90 0.78 0.95 0.97 0.44
Split-EN-10 0.76 0.46 0.91 0.82 1.00 0.89 0.73 0.95 0.96 0.51 0.91 0.79 0.96 0.97 0.42
Lasso 0.73 0.36 0.91 0.76 1.11 0.83 0.60 0.93 0.90 0.77 0.84 0.62 0.93 0.91 0.73
Elastic Net 0.74 0.38 0.91 0.78 1.08 0.85 0.66 0.94 0.93 0.66 0.87 0.70 0.94 0.94 0.60
Adaptive 0.68 0.06 0.98 0.60 1.25 0.77 0.32 0.96 0.83 0.96 0.78 0.34 0.96 0.84 0.93
Relaxed 0.72 0.34 0.90 0.75 1.17 0.82 0.63 0.91 0.90 0.94 0.83 0.63 0.91 0.90 1.27
0.3 MCP 0.70 0.26 0.91 0.69 1.23 0.76 0.44 0.90 0.80 1.05 0.77 0.46 0.90 0.81 1.01
SIS-SCAD 0.69 0.31 0.88 0.70 1.26 0.74 0.42 0.89 0.78 1.20 0.75 0.46 0.87 0.79 1.09
RuleFit 0.71 0.41 0.86 0.73 1.58 0.76 0.44 0.90 0.80 1.20 0.76 0.50 0.88 0.79 1.35
RE-Lasso-100 0.77 0.49 0.91 0.83 1.01 0.87 0.65 0.96 0.96 0.57 0.89 0.70 0.97 0.97 0.51
RE-EN-100 0.77 0.50 0.90 0.83 1.01 0.88 0.69 0.96 0.96 0.53 0.90 0.73 0.96 0.97 0.47
RGLM-100 0.74 0.27 0.96 0.82 1.06 0.81 0.40 0.99 0.96 0.85 0.83 0.45 0.99 0.97 0.83
RF-500 0.77 0.50 0.90 0.82 0.97 0.85 0.56 0.98 0.95 0.73 0.87 0.62 0.98 0.96 0.70
XGB 0.70 0.41 0.84 0.70 1.19 0.77 0.45 0.91 0.81 1.00 0.78 0.50 0.89 0.83 0.95
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.79 0.29 0.95 0.81 0.92 0.89 0.54 0.98 0.96 0.47 0.91 0.62 0.98 0.97 0.40
Split-EN-10 0.79 0.29 0.96 0.82 0.91 0.90 0.57 0.98 0.97 0.45 0.92 0.64 0.98 0.98 0.38
Lasso 0.78 0.24 0.95 0.75 1.04 0.85 0.40 0.97 0.90 0.68 0.86 0.43 0.97 0.91 0.64
Elastic Net 0.78 0.26 0.95 0.78 0.98 0.87 0.45 0.97 0.93 0.59 0.88 0.50 0.98 0.94 0.54
Adaptive 0.76 0.03 1.00 0.60 1.09 0.82 0.10 0.99 0.74 0.87 0.82 0.09 0.99 0.77 0.86
Relaxed 0.77 0.24 0.94 0.73 1.32 0.85 0.46 0.95 0.89 1.03 0.85 0.49 0.94 0.90 0.97
0.2 MCP 0.76 0.17 0.94 0.68 1.17 0.81 0.20 0.96 0.75 0.93 0.81 0.25 0.95 0.80 0.87
SIS-SCAD 0.75 0.16 0.94 0.70 1.17 0.81 0.17 0.96 0.77 0.93 0.81 0.23 0.95 0.78 0.88
RuleFit 0.76 0.27 0.91 0.69 1.50 0.80 0.30 0.93 0.74 1.28 0.80 0.35 0.92 0.75 1.35
RE-Lasso-100 0.80 0.35 0.94 0.83 0.91 0.89 0.51 0.98 0.96 0.48 0.90 0.54 0.99 0.97 0.43
RE-EN-100 0.80 0.35 0.94 0.83 0.91 0.90 0.56 0.98 0.96 0.44 0.91 0.57 0.99 0.97 0.40
RGLM-100 0.78 0.11 0.99 0.81 0.95 0.83 0.15 1.00 0.96 0.70 0.84 0.18 1.00 0.97 0.68
RF-500 0.80 0.31 0.95 0.82 0.87 0.86 0.32 0.99 0.95 0.61 0.87 0.37 1.00 0.96 0.59
XGB 0.76 0.26 0.91 0.71 1.07 0.82 0.29 0.95 0.82 0.83 0.83 0.32 0.95 0.83 0.80
Table 14: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.5, $n=$ 50, $p=$ 1000.
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.83 0.76 0.88 0.92 0.76 0.92 0.91 0.93 0.98 0.35 0.94 0.91 0.96 0.99 0.30
Split-EN-10 0.83 0.76 0.88 0.92 0.76 0.93 0.91 0.93 0.99 0.35 0.94 0.91 0.96 0.99 0.29
Lasso 0.81 0.73 0.87 0.90 0.85 0.88 0.86 0.90 0.96 0.52 0.90 0.85 0.93 0.97 0.47
Elastic Net 0.82 0.75 0.87 0.91 0.81 0.90 0.88 0.92 0.97 0.43 0.92 0.88 0.94 0.98 0.38
Adaptive 0.76 0.55 0.91 0.87 0.99 0.85 0.75 0.92 0.95 0.69 0.87 0.79 0.94 0.96 0.62
Relaxed 0.81 0.73 0.86 0.90 1.08 0.88 0.86 0.89 0.96 0.77 0.89 0.86 0.91 0.97 0.72
0.4 MCP 0.77 0.68 0.84 0.86 0.98 0.82 0.76 0.86 0.91 0.80 0.84 0.77 0.89 0.93 0.72
SIS-SCAD 0.76 0.64 0.85 0.85 1.05 0.79 0.70 0.85 0.89 0.89 0.81 0.71 0.88 0.90 0.82
RuleFit 0.77 0.68 0.84 0.84 1.33 0.80 0.72 0.86 0.88 1.17 0.81 0.71 0.88 0.90 1.04
RE-Lasso-100 0.83 0.77 0.88 0.92 0.76 0.92 0.89 0.93 0.98 0.39 0.93 0.89 0.96 0.99 0.34
RE-EN-100 0.83 0.77 0.88 0.92 0.76 0.92 0.90 0.93 0.98 0.37 0.94 0.90 0.96 0.99 0.32
RGLM-100 0.83 0.72 0.91 0.92 0.83 0.91 0.85 0.95 0.98 0.65 0.92 0.86 0.97 0.99 0.60
RF-500 0.83 0.76 0.89 0.92 0.74 0.91 0.87 0.94 0.98 0.56 0.92 0.87 0.96 0.99 0.53
XGB 0.78 0.70 0.84 0.87 0.95 0.83 0.77 0.87 0.92 0.77 0.84 0.75 0.90 0.93 0.73
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.84 0.69 0.92 0.92 0.72 0.93 0.83 0.97 0.99 0.33 0.94 0.88 0.97 0.99 0.28
Split-EN-10 0.85 0.69 0.92 0.92 0.71 0.93 0.84 0.97 0.99 0.32 0.94 0.89 0.97 0.99 0.27
Lasso 0.82 0.64 0.92 0.90 0.82 0.89 0.75 0.95 0.97 0.50 0.90 0.79 0.95 0.97 0.45
Elastic Net 0.83 0.66 0.92 0.91 0.77 0.91 0.79 0.96 0.98 0.42 0.92 0.84 0.96 0.98 0.35
Adaptive 0.77 0.37 0.96 0.89 0.93 0.84 0.54 0.98 0.95 0.68 0.86 0.63 0.97 0.96 0.61
Relaxed 0.82 0.65 0.91 0.90 0.99 0.88 0.77 0.93 0.96 0.96 0.89 0.79 0.93 0.96 0.95
0.3 MCP 0.78 0.52 0.91 0.85 0.96 0.82 0.57 0.93 0.89 0.84 0.83 0.61 0.92 0.90 0.82
SIS-SCAD 0.77 0.49 0.91 0.85 0.97 0.81 0.52 0.93 0.89 0.83 0.81 0.55 0.93 0.89 0.82
RuleFit 0.79 0.57 0.90 0.85 1.15 0.83 0.57 0.94 0.88 1.00 0.83 0.60 0.93 0.90 0.88
RE-Lasso-100 0.85 0.70 0.92 0.92 0.72 0.93 0.82 0.97 0.99 0.34 0.93 0.84 0.97 0.99 0.32
RE-EN-100 0.85 0.71 0.92 0.93 0.72 0.93 0.83 0.97 0.99 0.33 0.94 0.87 0.97 0.99 0.30
RGLM-100 0.84 0.61 0.95 0.92 0.77 0.90 0.69 0.99 0.99 0.58 0.91 0.73 0.99 0.99 0.56
RF-500 0.84 0.68 0.93 0.92 0.70 0.91 0.75 0.98 0.98 0.51 0.92 0.79 0.98 0.99 0.49
XGB 0.79 0.59 0.90 0.87 0.91 0.84 0.63 0.94 0.92 0.74 0.85 0.65 0.94 0.92 0.71
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.86 0.60 0.94 0.93 0.63 0.95 0.79 0.98 0.99 0.25 0.95 0.85 0.98 0.99 0.23
Split-EN-10 0.86 0.61 0.94 0.93 0.63 0.95 0.81 0.98 0.99 0.24 0.95 0.86 0.98 0.99 0.22
Lasso 0.84 0.53 0.94 0.91 0.72 0.91 0.66 0.97 0.97 0.42 0.92 0.74 0.96 0.97 0.38
Elastic Net 0.85 0.56 0.94 0.92 0.68 0.93 0.72 0.98 0.98 0.33 0.93 0.79 0.97 0.98 0.30
Adaptive 0.80 0.20 0.98 0.84 0.88 0.87 0.37 0.99 0.95 0.57 0.88 0.46 0.99 0.96 0.54
Relaxed 0.84 0.53 0.94 0.90 1.24 0.90 0.70 0.95 0.96 0.74 0.91 0.76 0.94 0.96 0.68
0.2 MCP 0.80 0.35 0.94 0.83 0.91 0.85 0.41 0.96 0.88 0.69 0.85 0.47 0.95 0.89 0.67
SIS-SCAD 0.79 0.31 0.94 0.84 0.95 0.84 0.32 0.97 0.89 0.72 0.85 0.41 0.95 0.89 0.73
RuleFit 0.81 0.47 0.91 0.80 1.34 0.86 0.47 0.95 0.85 0.99 0.86 0.52 0.95 0.87 0.97
RE-Lasso-100 0.87 0.66 0.93 0.93 0.62 0.94 0.78 0.98 0.99 0.27 0.95 0.83 0.98 0.99 0.26
RE-EN-100 0.87 0.66 0.93 0.93 0.62 0.95 0.81 0.98 0.99 0.25 0.95 0.84 0.98 0.99 0.25
RGLM-100 0.86 0.51 0.97 0.93 0.66 0.91 0.55 1.00 0.99 0.46 0.92 0.61 1.00 0.99 0.47
RF-500 0.87 0.61 0.95 0.93 0.61 0.92 0.63 0.99 0.99 0.41 0.93 0.71 0.99 0.99 0.41
XGB 0.82 0.49 0.92 0.86 0.84 0.87 0.51 0.95 0.91 0.66 0.87 0.56 0.94 0.92 0.66
Table 15: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.8, $n=$ 50, $p=$ 1000.
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.88 0.85 0.90 0.96 0.55 0.95 0.94 0.96 1.00 0.22 0.96 0.94 0.98 1.00 0.19
Split-EN-10 0.88 0.85 0.90 0.96 0.55 0.96 0.94 0.96 1.00 0.22 0.97 0.95 0.98 1.00 0.18
Lasso 0.87 0.84 0.89 0.95 0.61 0.93 0.90 0.95 0.99 0.32 0.94 0.91 0.96 0.99 0.28
Elastic Net 0.88 0.84 0.90 0.96 0.58 0.95 0.92 0.96 0.99 0.26 0.96 0.93 0.98 1.00 0.22
Adaptive 0.85 0.78 0.91 0.95 0.71 0.93 0.88 0.95 0.99 0.42 0.93 0.88 0.97 0.99 0.38
Relaxed 0.86 0.83 0.89 0.95 0.94 0.92 0.90 0.94 0.98 0.58 0.93 0.90 0.94 0.98 0.64
0.4 MCP 0.83 0.78 0.87 0.92 0.73 0.87 0.81 0.91 0.95 0.60 0.87 0.80 0.92 0.95 0.60
SIS-SCAD 0.84 0.77 0.88 0.93 0.74 0.88 0.82 0.92 0.96 0.59 0.88 0.80 0.94 0.96 0.59
RuleFit 0.84 0.80 0.88 0.91 1.01 0.88 0.83 0.91 0.91 1.01 0.88 0.83 0.91 0.90 1.17
RE-Lasso-100 0.88 0.85 0.90 0.96 0.56 0.95 0.93 0.96 0.99 0.25 0.96 0.94 0.98 1.00 0.22
RE-EN-100 0.88 0.86 0.90 0.96 0.56 0.95 0.94 0.96 1.00 0.23 0.96 0.94 0.98 1.00 0.21
RGLM-100 0.88 0.85 0.91 0.96 0.57 0.95 0.93 0.97 1.00 0.38 0.96 0.92 0.99 1.00 0.36
RF-500 0.88 0.85 0.90 0.96 0.54 0.95 0.93 0.96 0.99 0.35 0.96 0.93 0.98 1.00 0.32
XGB 0.84 0.80 0.87 0.91 0.80 0.87 0.83 0.90 0.90 0.75 0.87 0.83 0.89 0.89 0.80
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.89 0.81 0.93 0.96 0.51 0.95 0.92 0.97 0.99 0.21 0.96 0.94 0.97 1.00 0.19
Split-EN-10 0.89 0.81 0.93 0.96 0.51 0.96 0.93 0.97 1.00 0.21 0.97 0.95 0.97 1.00 0.18
Lasso 0.88 0.78 0.93 0.95 0.59 0.93 0.87 0.96 0.99 0.31 0.94 0.89 0.96 0.99 0.28
Elastic Net 0.89 0.80 0.93 0.96 0.54 0.95 0.90 0.97 0.99 0.25 0.95 0.91 0.97 0.99 0.22
Adaptive 0.85 0.62 0.96 0.95 0.69 0.92 0.80 0.97 0.98 0.41 0.93 0.82 0.97 0.99 0.41
Relaxed 0.88 0.78 0.92 0.95 0.86 0.92 0.88 0.94 0.98 0.53 0.93 0.89 0.94 0.98 0.77
0.3 MCP 0.84 0.69 0.91 0.92 0.71 0.87 0.72 0.93 0.94 0.60 0.87 0.74 0.93 0.95 0.59
SIS-SCAD 0.84 0.63 0.93 0.93 0.74 0.87 0.68 0.95 0.95 0.62 0.88 0.73 0.95 0.96 0.59
RuleFit 0.87 0.75 0.92 0.93 0.77 0.87 0.76 0.92 0.87 1.29 0.88 0.79 0.92 0.88 1.21
RE-Lasso-100 0.89 0.82 0.93 0.96 0.50 0.95 0.91 0.97 0.99 0.24 0.96 0.93 0.98 1.00 0.22
RE-EN-100 0.89 0.82 0.93 0.96 0.50 0.95 0.92 0.97 1.00 0.23 0.96 0.93 0.98 1.00 0.21
RGLM-100 0.89 0.80 0.94 0.96 0.52 0.95 0.88 0.98 1.00 0.35 0.96 0.90 0.99 1.00 0.35
RF-500 0.89 0.81 0.93 0.96 0.50 0.95 0.90 0.97 0.99 0.32 0.96 0.91 0.98 1.00 0.32
XGB 0.86 0.76 0.91 0.92 0.70 0.87 0.76 0.91 0.90 0.77 0.87 0.80 0.91 0.89 0.77
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.91 0.74 0.96 0.97 0.42 0.96 0.88 0.98 1.00 0.18 0.97 0.91 0.99 1.00 0.15
Split-EN-10 0.91 0.74 0.96 0.97 0.42 0.96 0.88 0.99 1.00 0.18 0.97 0.91 0.99 1.00 0.15
Lasso 0.90 0.69 0.96 0.96 0.49 0.94 0.80 0.98 0.99 0.28 0.95 0.82 0.98 0.99 0.24
Elastic Net 0.91 0.72 0.97 0.96 0.45 0.95 0.83 0.98 0.99 0.22 0.96 0.86 0.99 1.00 0.18
Adaptive 0.86 0.42 0.99 0.93 0.63 0.91 0.63 0.99 0.98 0.38 0.92 0.66 0.99 0.99 0.36
Relaxed 0.90 0.70 0.95 0.95 0.84 0.93 0.81 0.96 0.98 0.68 0.94 0.83 0.97 0.98 0.62
0.2 MCP 0.86 0.52 0.96 0.92 0.63 0.88 0.60 0.96 0.94 0.52 0.89 0.62 0.96 0.95 0.49
SIS-SCAD 0.85 0.44 0.97 0.93 0.65 0.88 0.54 0.97 0.95 0.54 0.89 0.55 0.98 0.96 0.51
RuleFit 0.87 0.66 0.94 0.85 1.19 0.88 0.71 0.93 0.82 1.30 0.90 0.71 0.94 0.83 1.16
RE-Lasso-100 0.91 0.76 0.96 0.97 0.41 0.96 0.87 0.98 1.00 0.20 0.97 0.91 0.99 1.00 0.17
RE-EN-100 0.91 0.77 0.96 0.97 0.41 0.96 0.88 0.98 1.00 0.19 0.98 0.92 0.99 1.00 0.15
RGLM-100 0.91 0.72 0.97 0.97 0.44 0.95 0.78 0.99 1.00 0.31 0.96 0.81 1.00 1.00 0.28
RF-500 0.91 0.74 0.97 0.97 0.43 0.95 0.81 0.99 0.99 0.29 0.96 0.83 0.99 1.00 0.26
XGB 0.87 0.68 0.93 0.91 0.69 0.88 0.71 0.93 0.93 0.65 0.90 0.71 0.94 0.94 0.58
Table 16: Mean prediction accuracy (ACC), sensitivity (SNS) and specificity (SPC), area under ROC curve (AUC) and test-sample loss (TSL) for Scenario 1, $\rho=$ 0.2, $n=$ 100, $p=$ 1000.
$\mathbf{\pi_{1}}$ Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 0.75 0.66 0.82 0.83 1.02 0.88 0.84 0.91 0.96 0.52 0.91 0.88 0.93 0.97 0.42
Split-EN-10 0.75 0.66 0.82 0.83 1.02 0.88 0.84 0.91 0.96 0.52 0.91 0.88 0.93 0.97 0.41
Lasso 0.72 0.61 0.80 0.80 1.10 0.84 0.77 0.89 0.93 0.71 0.86 0.81 0.90 0.94 0.64
Elastic Net 0.73 0.62 0.81 0.80 1.09 0.85 0.79 0.89 0.94 0.65 0.88 0.83 0.91 0.95 0.57
Adaptive 0.67 0.39 0.88 0.74 1.21 0.82 0.69 0.90 0.91 0.83 0.84 0.74 0.91 0.93 0.76
Relaxed 0.72 0.61 0.80 0.80 1.12 0.84 0.77 0.88 0.92 0.76 0.86 0.81 0.89 0.94 0.71
0.4 MCP 0.69 0.55 0.79 0.75 1.18 0.79 0.68 0.86 0.87 0.90 0.79 0.71 0.85 0.88 0.90
SIS-SCAD 0.68 0.56 0.77 0.74 1.21 0.75 0.65 0.82 0.82 1.06 0.75 0.66 0.82 0.83 1.05
RuleFit 0.71 0.61 0.78 0.77 1.64 0.79 0.68 0.87 0.88 1.01 0.80 0.69 0.87 0.88 0.97
RE-Lasso-100 0.75 0.67 0.81 0.83 1.04 0.88 0.82 0.92 0.96 0.54 0.90 0.85 0.94 0.97 0.45
[…]